Calculate the pH of 50 ml of 0.1000 M methylamine solution after the following volume of 0.1000 M HNO3, where added 0

1
Expert's answer
2022-06-14T20:25:04-0400

n(CH3NH2) = CM * V = 0.1 * 0.05 = 0.005mol

n(HNO3) = CM * V = 0.1 * 0 = 0 mol


CH3NH2 + H2O = CH3NH3+ + OH-

1mol CH3NH2 => 1mol OH-

0.1 M => X = 0.1 M OH-


Methylamine, CH3NH2, is a weak base that reacts with water according to the above equation. For this reaction at equilibrium, Kb = 4.4 x 10-4

CM (OH-) = 0.1 * 4.4*10-4 = 4.4*10-5 M


pOH = -log [OH-] = - log (4.4*10-5) = 4.357


Using pH + pOH = 14, we find the pH value:

pH = 14 - pOH = 14 - 4.357 = 9.643


Answer: pH = 9.643